OpenStudy (anonymous):

Help Please! Find the value of the constant c that makes the following function continuous on (-infinity, +infinity) f(x)={ x^2-c if -infinity

OpenStudy (anonymous):

@tkhunny

OpenStudy (helder_edwin):

u have \[\large f(x)=\begin{cases} x^2-c, &x<5\\ cx+5, & x\geq 5 \end{cases} \]

OpenStudy (helder_edwin):

use the definition of continuity \[\large \lim_{x\to5}f(x)=f(5) \]

OpenStudy (helder_edwin):

and the existence condition for limits: both lateral limits must exist and \[\large \lim_{x\to5+}f(x)=\lim_{x\to5-}f(x) \]

OpenStudy (anonymous):

\[f(x)=\left[\begin{matrix}x^2-c, & if -\infty<x<5 \\ cx+5, & if x \ge5\end{matrix}\right]\] I understand what you've explained but how would I find c.?

OpenStudy (anonymous):

@helder_edwin

OpenStudy (anonymous):

@Loser66 help?(:

OpenStudy (helder_edwin):

ok. for the left limit, since \(x\to5-\) then \(x<5\) so \[\large \lim_{x\to5-}f(x)=\lim_{x\to5-}(x^2-c)=5^2-c=25-c \] for the right limit, since \(x\to5+\) then \(x>5\) so \[\large \lim_{x\to5+}f(x)=\lim_{x\to5+}(cx+5)=5c+5. \]

OpenStudy (helder_edwin):

now, for the function to be continous, these limits have to be equal so \[\large 25-c=5c+5 \] \[\large 20=6c \] \[\large c=\frac{20}{6}=\frac{10}{3} \]

OpenStudy (anonymous):

oh! that made so much sense, I kinda over thought it. Thank you!!! @helder_edwin

OpenStudy (helder_edwin):

u r welcome